OpenStudy (anonymous):

-4(x + 1) - 3 = -3(x - 4) -19 19 2 0

OpenStudy (jbo11):

distribute the -4 and -3 to get rid of the ( ) and then solve for x

OpenStudy (anonymous):

just give me the steps and awnser. @Jbo11

OpenStudy (jbo11):

once you distribute you get -4x-4-3=-3x+12 Then get the x's to one side so -4x+3x-4-3=12 then the other numbers to the other side so -4x-3x=12+4+3 then solve for x

OpenStudy (anonymous):

so it would be 19?@Jbo11

OpenStudy (anonymous):

it says it would be -19

OpenStudy (jbo11):

yes it's -19 -4+3x = -1x or -x so then it's -x=12+4+3 which is 19 then you need to divide the negative to make x positive so x= 19/-1 which is -19

OpenStudy (anonymous):

can u hewlp me with this assignment?@Jbo11

OpenStudy (jbo11):

haha well, there are many others on this website that can help too, just post your questions and people will help, I'll try but I may have to go somewhere soon